# POPoS: Improving Efficient and Robust Facial Landmark Detection with Parallel Optimal Position Search

### v1

@@ Abstract
Achieving a balance between accuracy and efficiency is a critical challenge in facial landmark detection (FLD). This paper introduces the Parallel Optimal Position Search (POPoS), a high-precision encoding-decoding framework designed to address the fundamental limitations of traditional FLD methods. POPoS employs three key innovations: (1) Pseudo-range multilateration is utilized to correct heatmap errors, enhancing the precision of landmark localization. By integrating multiple anchor points, this approach minimizes the impact of individual heatmap inaccuracies, leading to robust overall positioning. (2) To improve the pseudo-range accuracy of selected anchor points, a new loss function, named multilateration anchor loss, is proposed. This loss function effectively enhances the accuracy of the distance map, mitigates the risk of local optima, and ensures optimal solutions. (3) A single-step parallel computation algorithm is introduced, significantly enhancing computational efficiency and reducing processing time. Comprehensive evaluations across five benchmark datasets demonstrate that POPoS consistently outperforms existing methods, particularly excelling in low-resolution scenarios with minimal computational overhead. These features establish POPoS as a highly efficient and accurate tool for FLD, with broad applicability in real-world scenarios. The code is available at https://github.com/teslatasy/PoPoS.
Introduction
@@
Facial Landmark Detection (FLD) is a fundamental cornerstone in the field of computer vision, playing an indispensable role in the accurate identification and localization of salient facial features such as eye regions, nasal structures, and mouth contours in digital images. This crucial task underpins a wide range of advanced applications, including sophisticated facial recognition systems (Cheng et al. 2019; Huang et al. 2020b; Cheng et al. 2022; Liu et al. 2023), intricate three-dimensional face modeling techniques (Tu et al. 2023; Zhou et al. 2023a), and critical advances in autonomous vehicle technologies (Kim et al. 2023; Ghoddoosian, Galib, and Athitsos 2019). The importance of FLD extends far beyond mere detection; it serves as the linchpin for interpreting and analyzing human facial expressions, emotional states, and identity markers in digital environments, catalyzing innovation in fields as diverse as human-computer interaction paradigms, augmented reality systems, and medical diagnostic protocols.
@@
The advent of deep learning methods has brought about a paradigm shift in FLD research, primarily through two dominant approaches: heatmap-based methods and coordinate regression techniques. Heatmap regression methods (Chandran et al. 2020; Zhang et al. 2020; Jin, Liao, and Shao 2021) have received considerable attention due to their ability to generate probabilistic heatmaps for individual landmarks, effectively preserving spatial and contextual information while mitigating the risk of overfitting during the training process. These methods are remarkably effective at capturing the inherent uncertainty associated with landmark positions, providing robust performance across a wide range of facial expressions and orientations.
@@
Figure 1: Visualization of quantization error in the coordinate decoding process. The discrepancy between the ground truth (GT) coordinates and the point of highest probability in the heatmap illustrates the challenge of achieving lossless decoding and precise localization.
@@
Despite their merits, heatmap-based approaches are not without challenges. One significant obstacle is the occurrence of quantization errors, particularly when downscaling heatmaps from their original resolution (Bao et al. 2023). This problem arises from the inevitable discretization of continuous spatial information into pixelated representations, resulting in a potential degradation of precision during the landmark localization process. While maintaining high-resolution heatmaps can partially mitigate this problem, it imposes significant computational requirements, making it impractical for real-time applications or deployment on resource-constrained devices.
@@
The heart of the challenge in FLD is therefore to achieve an optimal balance between high-resolution accuracy and computational efficiency. This delicate balance becomes particularly critical in scenarios requiring real-time performance or in environments characterized by limited processing power. Although seminal works such as Hourglass (Newell, Yang, and Deng 2016), Crossnet (Zhang et al. 2022), and KeyposS (Bao et al. 2023) have made significant progress in addressing accuracy degradation in low-resolution heatmaps, the pursuit of an optimal trade-off between accuracy and efficiency remains an open challenge in the field. State-of-the-art methods often struggle to maintain high accuracy when faced with low-resolution heatmaps or stringent computational constraints.
@@
This persistent challenge underscores the imperative for innovative encoding and decoding techniques that can overcome the inherent limitations of current paradigms. The research community is actively exploring novel approaches that leverage the strengths of heatmap-based methods while addressing their shortcomings, potentially through the development of hybrid architectures or advanced post-processing strategies. Fig. 1 illustrates a fundamental challenge in FLD: the quantization error inherent in the coordinate decoding process. This visual representation delineates the potential misalignment between the ground truth (GT) landmark coordinates and the point of highest probability as indicated by the heatmap. This discrepancy is a direct consequence of the discretization process, wherein the continuous spatial domain is mapped onto a discrete pixel grid, potentially compromising localization accuracy. The figure serves as a compelling argument for the necessity of sophisticated decoding strategies capable of bridging the gap between discretized heatmap representations and the continuous spatial domain of facial features.
@@
To address these multifaceted challenges, we propose the Parallel Optimal Position Search (POPoS) framework, a novel approach designed to overcome the accuracy-efficiency trade-offs inherent in heatmap-based methods. POPoS introduces a series of innovative strategies that collectively address the core limitations of traditional FLD approaches. At its core, POPoS employs a pseudo-range multilateration technique that uses the top K points of the heatmap as anchors, effectively mitigating the impact of multiple peaks and ensuring accurate landmark localization. This is complemented by a specialized anchor-based optimization strategy that balances global and local optimization, compensating for quantization errors and improving sub-pixel accuracy. Furthermore, POPoS implements a single-step parallel computation algorithm that dramatically reduces computation time while maintaining high accuracy, making it particularly well-suited for modern GPU architectures. These improvements enable POPoS to demonstrate particular efficacy in low-resolution heatmaps scenarios without compromising on accuracy or computational efficiency, achieving an optimal balance between precision and efficiency. This positions POPoS as a viable solution for a wide spectrum of real-world applications, from mobile devices to advanced real-time systems. The key contributions of POPoS are:
@@
  • A novel pseudo-range multilateration approach that outperforms traditional methods in scenarios with multiple local maxima or ambiguous heatmap regions.
@@
  • Extensive evaluation across multiple benchmarks, demonstrating superior performance in low-resolution scenarios with minimal computational overhead.
Related Work
Facial landmark recognition has advanced through two key approaches: heatmap regression and high-efficiency heatmap decoding.
Heatmap Regression-Based Methods
Heatmap regression methods (Chandran et al. 2020; Sun et al. 2019; Tang et al. 2018; Zou et al. 2019; Wang, Bo, and Li 2019; Jin, Liao, and Shao 2021) have evolved to predict landmarks using high-resolution feature maps with increasing sophistication. The Stacked Hourglass Network (Newell, Yang, and Deng 2016) and UNet (Ronneberger, Fischer, and Brox 2015) established the foundation by preserving spatial relationships. Subsequent innovations addressed specific challenges: UDP (Huang et al. 2020a) tackled discretization errors, G-RMI (Papandreou et al. 2017) mitigated quantization errors by combining dense heatmaps with keypoint offsets, and the DSNT layer (Nibali et al. 2018) introduced coordinate supervision regularization. PIPNet (Jin, Liao, and Shao 2021) further streamlined the process with simultaneous predictions on low-resolution feature maps. This progression has consistently enhanced accuracy by leveraging spatial pixel relationships.
High Efficiency Heatmap Decoding Methods
While heatmap regression improved prediction accuracy, extracting coordinates from estimated heatmaps remained a critical area for optimization. High-efficiency decoding methods emerged to address this challenge. Darkpose (Zhang et al. 2020) refined heatmap generation and distribution analysis, while FHR (Tai et al. 2019) enhanced precision through improved fractional part estimation. Subpixel heatmap regression (Bulat, Sanchez, and Tzimiropoulos 2021) increased accuracy by predicting offsets using soft continuous maximum values. KeyPosS (Bao et al. 2023) introduced a novel approach by integrating GPS-inspired true-range multi-point positioning technology.
Building on these advancements, our Parallel Optimal Position Search (POPoS) method combines pseudo-range multilateration with Potential Position Parallel Sampling. POPoS addresses limitations in predictive accuracy and computational efficiency, particularly at lower heatmap resolutions. This innovation not only advances facial landmark detection but also extends its applicability to resource-constrained environments, marking a significant step towards more accessible and efficient high-accuracy detection across diverse applications and hardware configurations.
Parallel Optimal Position Search (POPoS)
@@
These interconnected challenges hinder robust, accurate, and efficient facial landmark detection across diverse scenarios. To address these challenges, we propose Parallel Optimal Position Search (POPoS), a novel framework that fundamentally reimagines heatmap-based facial landmark detection. POPoS comprises three stages, as illustrated in Fig. 2. First, the model optimization phase involves the generation of heatmaps and optimization, where the multilateration anchor loss is conducted to reduce encoding information loss. Second, the heatmap generation and anchor sampling phase include model inference, distance transformation, and anchor point selection. Third, the sample and search phase involves constructing the pseudo-range positioning equation, parallel sampling, and equivalent single-step parallel computation.
Figure 2: Parallel Optimal Position Search (POPoS) framework. The Multilateration Anchor Loss is designed during the training phase to optimize distance prediction. During model inference, POPoS samples the top-K response positions as anchor points for distance map decoding.
Model & Optimization
@@
where $\mathcal{H}_{i}$ and ^$\hat{\mathcal{H}}_{i}$ are the ground truth and predicted heatmaps, and $\text{Mask}_{i}\in R^{h\times w}$ indicates a binary mask which assigns the top K response locations with value one, while the others remain zero, $i$ represents the $i$-th landmark.
Further, the model is jointly optimized using MSE loss and MA loss. The loss $L$ in the training stage is $\mathcal{L}=\mathcal{L}_{\text{MSE}}+{w}\times\mathcal{L}_{\text{MA}}$, where $\mathcal{L_{\text{MSE}}}$ is the MSE loss of the whole heatmap, $\mathcal{L}_{\text{MA}}$ is the MA loss, and ${w}$ is the weight parameter of local anchor loss.
Heatmap Generation & Anchor Sampling
In this stage, there are three key steps: heatmap generation, distance transformation, and anchor sampling.
First step is heatmap generation: An input image $I\in\mathbb{R}^{H\times W\times 3}$ is processed through a streamlined HRNet variant (Bao et al. 2023) to extract features. The heatmaps $\mathcal{H}\in\mathbb{R}^{h\times w\times N_{k}}$ is then generated, where $h=H/\lambda$, $w=W/\lambda$, and $N_{k}$ represents the number of facial landmarks. The unbiased coordination encoding is employed during the encoding stage. The coordinate heatmap is generated as a 2-dimensional Gaussian distribution/kernel centered at the labeled coordinate.
@@
The second step is anchor sampling. To mitigate the influence of prediction errors induced by heatmap generation, different from the triangulation algorithm (Tai et al. 2019), the top K points of the distance map are selected as anchors. Because the selected top K points are consistent with those top K points in the model optimization stage, they promise to minimize information loss. Meanwhile, these selected top K points are utilized to form distance equations in the next stage.
Parallel Sampling & Computing
This stage involve two steps, forming pseudo-range multilateration equation, potential position parallel sampling and computing.
@@
where $\varepsilon_{i}$ represents the error between the decoded distance and the actual distance.
A novel method for solving the objective function is proposed, called Potential Position Parallel Sampling and Computing (PPPSC). Through three steps (potential region selection, candidate generation, and Solution of Parallel Computing), it can directly find the point that minimizes the objective function.
Potential Region selection. First, considering the subpixel scenario, there are an infinite number of points existing in the heatmap. Second, as the distance between the sample point with the maximum value point $m$ in the heatmap increases, the probability of samping point hitting the point of interest decreases. Third, an excess of sampling can lead to redundant computations. Hence, the search area is set as a square region of $h\times w$ centered at point $m$.
Candidates Generation. Within the selected region, sampling is conducted using a mesh with a frequency of $\tau$ samples per pixel and obtained a total of $N$ points. Formally, a mesh grid point set $G=\{g_{1},\dots,g_{N}\}$ is generated first, $g_{i}=(\frac{\Delta\bm{\hat{}}{u}_{i}}{\lambda},\frac{\Delta\bm{\hat{}}{v}_{i}}{\lambda})$, where $\Delta\bm{\hat{}}{u}_{i}\in[0,\lambda\times h\times\tau],\Delta\bm{\hat{}}{v}_{i}\in[0,\lambda\times w\times\tau]$. And $\tau$ is a factor to control the granularity of subpixel sampling. $\lambda$ is the downsampling factor. Then, the mesh grid point set $G$ is translated to cover the maximum value position $(\hat{u}_{m},\hat{v}_{m})~{}$in the grid center, and the candidates $c_{i}\in C$ is computed as Eq. 4:
@@
Solution of Parallel Computing.
After the sampling, the computing step for one heatmap is formed as Eq. 5. All points from the sampling set $C$ are then inputted into the objective function $\Phi(\mathcal{A},C)$, which calculates the discrepancy between two distance matrices: one is the Euclidean distance matrix between the sampled points $C$ and the anchor set $\mathcal{A}$, and the other one is the predicted distance map $D$ of anchor set $\mathcal{A}$. This results in an error matrix $E$. The optimal facial landmark position is determined by locating the minimum position index in the matrix $E$.
@@
Experiments
Dataset and Metric

### v5

@@ Abstract
Achieving a balance between accuracy and efficiency is a critical challenge in facial landmark detection (FLD). This paper introduces Parallel Optimal Position Search (POPoS), a high-precision encoding-decoding framework designed to address the limitations of traditional FLD methods. POPoS employs three key contributions: (1) Pseudo-range multilateration is utilized to correct heatmap errors, improving landmark localization accuracy. By integrating multiple anchor points, it reduces the impact of individual heatmap inaccuracies, leading to robust overall positioning. (2) To enhance the pseudo-range accuracy of selected anchor points, a new loss function, named multilateration anchor loss, is proposed. This loss function enhances the accuracy of the distance map, mitigates the risk of local optima, and ensures optimal solutions. (3) A single-step parallel computation algorithm is introduced, boosting computational efficiency and reducing processing time. Extensive evaluations across five benchmark datasets demonstrate that POPoS consistently outperforms existing methods, particularly excelling in low-resolution heatmaps scenarios with minimal computational overhead. These advantages make POPoS as a highly efficient and accurate tool for FLD, with broad applicability in real-world scenarios. The code is available at https://github.com/teslatasy/POPoS.
1 Introduction
@@
Facial Landmark Detection (FLD) is a fundamental task in computer vision, essential for identifying and localizing key facial features such as the eyes, nose, and mouth in digital images (Wu and Ji 2018). This crucial process underpins a wide array of advanced applications, including facial recognition systems (Khabarlak and Koriashkina 2021) and three-dimensional face modeling (Zeng et al. 2023). Beyond mere detection, FLD serves as a cornerstone for interpreting human facial expressions (Cheng et al. 2024b, c), analyzing emotional states (Cheng et al. 2024a), and facilitating interactions within digital environments (Cheng et al. 2016, 2017a, 2017b, 2017c). Consequently, FLD drives innovation across diverse fields such as human-computer interaction (Li et al. 2024; He et al. 2023, 2024, 2023) and augmented reality systems (Xu et al. 2024a; Tu et al. 2024a, b; Xu et al. 2024b).
@@
The advent of deep learning has advanced FLD research, primarily through two dominant approaches: heatmap-based methods and coordinate regression techniques. Heatmap regression methods (Chandran et al. 2020; Zhang et al. 2020; Jin, Liao, and Shao 2021) have garnered considerable attention due to their ability to generate probabilistic heatmaps for individual landmarks. These methods effectively preserve spatial and contextual information while mitigating the risk of overfitting during training. Additionally, they adeptly capture the inherent uncertainty in landmark positions, enabling robust performance across a wide range of facial expressions and orientations.
@@
Figure 1: Quantization error in the coordinate decoding process. The discrepancy between the ground truth (GT) coordinates and the highest probability point in the heatmap illustrates the challenge of achieving precise localization and lossless decoding. [Best viewed in color with zoom]
@@
Despite their strengths, heatmap-based approaches encounter significant challenges. A primary limitation is quantization error, which becomes pronounced when downscaling heatmaps from their original resolution (Bao et al. 2023). This issue stems from the discretization of continuous spatial information into pixelated representations, leading to potential precision loss during landmark localization. Although maintaining high-resolution heatmaps can mitigate this problem, it imposes substantial computational demands, rendering it impractical for real-time applications or deployment on resource-constrained devices.
@@
Furthermore, the core challenge in FLD lies in balancing localization accuracy with computational efficiency, particularly in scenarios that require real-time performance or operate under limited processing power. This trade-off is increasingly critical as applications demand both high precision and speed . While pioneering works such as Hourglass (Newell, Yang, and Deng 2016), CrossNet (Zhang et al. 2022), and KeyposS (Bao et al. 2023) have made notable strides in reducing accuracy degradation in low-resolution heatmaps, achieving an optimal balance between accuracy and efficiency remains an open challenge. Current state-of-the-art methods often struggle to maintain high accuracy under low-resolution conditions or stringent computational constraints, indicating a need for further exploration.
@@
This ongoing pursuit for balance in FLD is compounded by the inherent limitations of current encoding and decoding techniques, which further restrict localization performance. Research efforts are actively exploring ways to enhance heatmap-based approaches by leveraging their strengths while mitigating their weaknesses, often through hybrid architectures or advanced post-processing strategies. Fig. 1 illustrates a key challenge in FLD: quantization error in coordinate decoding. The figure highlights the misalignment between the ground truth (GT) landmark coordinates and the heatmap’s highest-probability point. This discrepancy arises from the discretization process, where the continuous spatial domain is mapped onto a discrete pixel grid, compromising localization accuracy. The visualization underscores the importance of advanced decoding strategies to bridge the gap between discrete heatmap representations and the continuous spatial domain of facial features. These limitations emphasize the pressing need for innovative methods to improve accuracy while addressing the challenges of discretization and quantization errors.
@@
To address these interconnected challenges, we present the Parallel Optimal Position Search (POPoS) framework, a novel solution designed to resolve the accuracy-efficiency trade-offs inherent in heatmap-based methods. POPoS introduces a set of complementary strategies to tackle the core limitations of traditional FLD approaches. At its foundation, POPoS employs a pseudo-range multilateration technique that leverages the top $K$ points of the heatmap as anchors, effectively reducing the influence of multiple peaks and enhancing landmark localization accuracy. This is further complemented by an anchor-based optimization strategy that integrates global and local refinements, compensating for quantization errors and achieving sub-pixel precision. In addition, POPoS incorporates a single-step parallel computation algorithm, significantly reducing computation time while preserving high accuracy—an approach well-suited to modern GPU architectures. These improvements enable POPoS to excel in low-resolution heatmap scenarios, striking an effective balance between precision and computational efficiency. Consequently, POPoS emerges as a versatile solution applicable to a wide range of real-world settings, from mobile platforms to real-time systems. The key contributions of POPoS are summarized as follows:
@@
  • A pseudo-range multilateration approach that improves localization in scenarios with multiple local maxima or ambiguous heatmap regions.
@@
  • Extensive evaluation across multiple datasets demonstrating superior performance in low-resolution heatmap scenarios, with minimal computational overhead.
2 Related Work
Facial landmark detection research has two key directions: heatmap regression and high-efficiency heatmap decoding.
2.1 Heatmap Regression-Based Methods
Heatmap regression methods (Chandran et al. 2020; Sun et al. 2019; Tang et al. 2018; Zou et al. 2019; Wang, Bo, and Li 2019; Jin, Liao, and Shao 2021; Cheng et al. 2019b, a, 2022; Huang et al. 2020a; Zhang et al. 2022) have progressed by predicting landmarks using high-resolution feature maps. The Stacked Hourglass Network (Newell, Yang, and Deng 2016) and UNet (Ronneberger, Fischer, and Brox 2015) established the foundation by preserving spatial relationships. Subsequent innovations addressed specific challenges: UDP (Huang et al. 2020b) tackled discretization errors, G-RMI (Papandreou et al. 2017) mitigated quantization errors by combining dense heatmaps with keypoint offsets, and the DSNT layer (Nibali et al. 2018) introduced coordinate supervision regularization. PIPNet (Jin, Liao, and Shao 2021) further streamlined the process with simultaneous predictions on low-resolution heatmaps. This progression has consistently enhanced accuracy by leveraging spatial pixel relationships (Zou et al. 2019; Wang, Bo, and Li 2019).
2.2 High Efficiency Heatmap Decoding Methods
While heatmap regression improved prediction accuracy, extracting coordinates from estimated heatmaps remained a critical area for optimization. High-efficiency decoding methods were developed to address this challenge. Darkpose (Zhang et al. 2020) refined heatmap generation and distribution analysis, while FHR (Tai et al. 2019) enhanced precision through improved fractional part estimation. Subpixel heatmap regression (Bulat, Sanchez, and Tzimiropoulos 2021) increased accuracy by predicting offsets using soft continuous maxima. KeyPosS (Bao et al. 2023) introduced a novel approach by integrating GPS-inspired multi-point positioning technology for true-range estimation.
Building on these advancements, our Parallel Optimal Position Search (POPoS) method combines pseudo-range multilateration and Potential Position Parallel Sampling. POPoS addresses limitations in predictive accuracy and computational efficiency, particularly in scenarios with low-resolution heatmaps. This innovation not only advances facial landmark detection but also extends its applicability to resource-constrained environments. It marks a step towards more accessible and efficient high-accuracy detection across diverse applications and hardware configurations.
3 Parallel Optimal Position Search (POPoS)
@@
These interconnected challenges hinder robust, accurate, and efficient facial landmark detection in a wide range of scenarios. To address these challenges, we propose Parallel Optimal Position Search (POPoS), a framework that fundamentally reimagines heatmap-based facial landmark detection. POPoS comprises three stages, as shown in Fig. 2. First, the model optimization phase involves generating heatmaps and optimizing them, where the multilateration anchor loss is applied to minimize encoding information loss. Second, the heatmap generation and anchor sampling phase include model inference, distance transformation, and anchor point selection. Third, the sample and search phase involves constructing the pseudo-range positioning equation, parallel sampling, and equivalent single-step parallel computation. Details are in the following sections.
Figure 2: Parallel Optimal Position Search (POPoS) framework. The Multilateration Anchor Loss is designed during the training phase to optimize distance prediction. During model inference, POPoS samples the top-K response positions as anchor points for distance map decoding. [Best viewed in color with zoom]
3.1 Model & Optimization
@@
where $\mathcal{H}_{i}$ and ^$\hat{\mathcal{H}}_{i}$ are the ground truth and predicted heatmaps, and $\text{Mask}_{i}\in R^{h\times w}$ indicates a binary mask which assigns the top K response locations with value one, while the others remain zero, $i$ represents the $i$-th landmark. Further, the model is jointly optimized using MSE loss and MA loss. The loss $L$ in the training stage is $\mathcal{L}=\mathcal{L}_{\text{MSE}}+{w}\times\mathcal{L}_{\text{MA}}$, where $\mathcal{L_{\text{MSE}}}$ is the MSE loss of the whole heatmap, $\mathcal{L}_{\text{MA}}$ is the MA loss, and ${w}$ is the weight parameter of local anchor loss.
3.2 Heatmap Generation & Anchor Sampling
In this stage, there are three key steps:1) heatmap generation, 2) distance transformation, and 3) anchor sampling.
First step is heatmap generation. An input image $I\in\mathbb{R}^{H\times W\times 3}$ is processed through a streamlined HRNet variant (Bao et al. 2023) to extract features. The heatmaps $\mathcal{H}\in\mathbb{R}^{h\times w\times N_{k}}$ is then generated, where $h=H/\lambda$, $w=W/\lambda$, and $N_{k}$ represents the number of facial landmarks. The unbiased coordination encoding is employed during the encoding stage. The coordinate heatmap is generated as a 2-dimensional Gaussian distribution/kernel centered at the labeled coordinate.
@@ 3.2 Heatmap Generation & Anchor Sampling
The third step is anchor sampling. To mitigate the prediction errors induced by heatmap generation, different from the triangulation algorithm (Tai et al. 2019), the top K points of the distance map are selected as anchors. Because the selected top K points are consistent with those top K points in the model optimization stage, they promise to minimize information loss. Meanwhile, these selected top K points are utilized to form distance equations in the next stage.
3.3 Parallel Sampling & Computing
This stage involve two key steps: 1) forming the pseudo-range multilateration equation, and 2) potential position parallel sampling and computing as follows:
@@ 3.3 Parallel Sampling & Computing
where $\varepsilon_{i}$ represents the error between the decoded distance and the actual distance. A new method for solving the objective function is proposed, called Potential Position Parallel Sampling and Computing (PPPSC). Through three steps (potential region selection, candidate generation, and Solution of Parallel Computing), it can directly find the point that minimizes the objective function. The detailed steps are outlined and further explained below.
(1) Potential Region Selection. First, considering the subpixel scenario, there are an infinite number of points existing in the heatmap. Second, as the distance between the sample point with the maximum value point $m$ in the heatmap increases, the probability of samping point hitting the point of interest decreases. Third, an excess of sampling can lead to redundant computations. Hence, the search area is set as a square region of $h\times w$ centered at point $m$.
(2) Candidates Generation. Within the selected region, sampling is conducted using a mesh with a frequency of $\tau$ samples per pixel and obtained a total of $N$ points. Formally, a mesh grid point set $G=\{g_{1},\dots,g_{N}\}$ is generated first, $g_{i}=(\frac{\Delta\bm{\hat{}}{u}_{i}}{\lambda},\frac{\Delta\bm{\hat{}}{v}_{i}}{\lambda})$, where $\Delta\bm{\hat{}}{u}_{i}\in[0,\lambda\times h\times\tau],\Delta\bm{\hat{}}{v}_{i}\in[0,\lambda\times w\times\tau]$. And $\tau$ is a factor to control the granularity of subpixel sampling. $\lambda$ is the downsampling factor. Then, the mesh grid point set $G$ is translated to cover the maximum value position $(\hat{u}_{m},\hat{v}_{m})~{}$in the grid center, and the candidates $c_{i}\in C$ is computed as Eq. 4:
@@
(3) Solution of Parallel Computing. After the sampling, the computing step for one heatmap is formed as Eq. 5. All points from the sampling set $C$ are then inputted into the objective function $\Phi(\mathcal{A},C)$, which calculates the discrepancy between two distance matrices: one is the Euclidean distance matrix between the sampled points $C$ and the anchor set $\mathcal{A}$, and the other one is the predicted distance map $D$ of anchor set $\mathcal{A}$. This results in an error matrix $E$. The optimal facial landmark position is determined by locating the minimum position index in the matrix $E$.
@@
4 Experiments
4.1 Dataset and Metric